# Multiply-Robust Causal Change Attribution

### v1

@@ Abstract
Comparing two samples of data, we observe a change in the distribution of an outcome variable. In the presence of multiple explanatory variables, how much of the change can be explained by each possible cause? We develop a new estimation strategy that, given a causal model, combines regression and re-weighting methods to quantify the contribution of each causal mechanism. Our proposed methodology is multiply robust, meaning that it still recovers the target parameter under partial misspecification. We prove that our estimator is consistent and asymptotically normal. Moreover, it can be incorporated into existing frameworks for causal attribution, such as Shapley values, which will inherit the consistency and large-sample distribution properties. Our method demonstrates excellent performance in Monte Carlo simulations, and we show its usefulness in an empirical application.
@@ 1 Introduction
Although other recent algorithms have been proposed to estimate Shapley Values, a majority of these algorithms are not multiply-robust, and generally they are based on regression only (learning the distribution of the outcome given the explanatory variables). [^1] The only exception that we are aware of is Jung et al. (2022). However, their notion of $do$-Shapley values decomposes the effect of fixing a value for the explanatory variables, rather than a distribution for each causal mechanism, as we do in this paper. They consider only discrete-valued covariates, and their approach does not seem easily generalizable beyond that. As such, they will not perform well unless we have a high-quality estimator of the regression function. In high-dimensional or non-parametric settings, estimators for those objects will typically exhibit slow convergence rates, rendering traditional normal-based large sample inference (standard errors, confidence intervals or $p$-values) invalid. We provide the first formal analysis of the asymptotic properties of an estimator in the causal change attribution setting of Budhathoki et al. (2021), which allows us to perform valid inference in large samples (standard errors, confidence intervals and tests).
Causal change attribution is related to the classical problem of treatment effect estimation (Pearl, 2009; Imbens & Rubin, 2015; Peters et al., 2017) in the need to evaluate counterfactual distributions, but there are some fundamental differences. Within the treatment effects literature, the closest problem to ours is that of causal mediation. Tchetgen-Tchetgen & Shpitser (2012) gave multiple-robustness results for mediation analysis with a single mediator (explanatory variable). The case with multiple mediators and an arbitrary DAG is substantially more challenging. Although causal mediation with multiple mediators has been studied before (Daniel et al., 2015), we provide the first multiply-robust estimator in this setting. We also believe to be the first to explicitly build a connection between the causal change attribution problem and the causal mediation literature.
@@ 2.1 Setting
We want to draw a clear distinction between attribution and a related problem that has received a much larger attention in the literature, Average Treatment Effect (ATE) estimation under unconfoundedness. The difference between the causal models underlying these two problems is depicted in Figure 1. In an ATE setting, $\bm{X}$ typically represents pre-treatment covariates, that affect both the outcome $Y$ and the propensity to receive a binary treatment $T$. To compute the causal effect of $T$ we need to control for those covariates, that is, compare units with similar values of $\bm{X}$. In contrast, causal change attribution acknowledges that both the distributions of $\bm{X}$ and of $Y$ given $\bm{X}$ vary depending on whether $T=0$ or $T=1$. The goal of causal change attribution is to quantify how much of the effect of $T$ on $Y$ goes through (is mediated by) each causal mechanism in this distribution. Under an unconfoundedness assumption about the assignment of $T$, the problem of causal attribution can be thought of as decomposing the ATE of $T$ on $Y$ into the Natural Direct Effect and the Natural Indirect Effect corresponding to each causal mechanism (Pearl, 2009, §4.5; Daniel et al., 2015). We describe this in more detail, including the corresponding structural equation model and potential outcomes, in Appendix B.
We note, however, that our definition of attribution does not require $T$ to be a treatment that can be administered in the “interventional” sense. As discussed, $T$ could also be an indicator for group membership (e.g., female or male), time period (e.g., before and after a certain date), or geographical location (e.g., East vs. West Coast of the US). Under which conditions can this change attribution be interpreted as causal? We will assume that the researcher knows a causal ordering of the underlying true causal graph for $(\bm{X},Y)$. This will allow us to quantify the contribution of each causal mechanism, rather than changes in the marginal distributions of covariates. In practice, researchers can use a combination of causal discovery methods (e.g., the conditional independence test of Zhang et al., 2011) and domain knowledge to obtain the DAG. We also assume that there is no unobserved variable $U$ such that $T\rightarrow U$, $U\rightarrow\bm{X}$ and $U\rightarrow Y$. We discuss ways to relax this assumption in future research in Section 4.
@@ 2.2 Preliminary Example
$\theta^{\langle 1,0\rangle}:=\int y\mathrm{d}P^{\langle 1,0\rangle}_{Y}(y),\quad\text{for}\quad P^{\langle 1,0\rangle}_{Y}:=\int P_{Y\mid X}^{(0)}\mathrm{d}P_{X}^{(1)}.$
The fundamental challenge is that $P^{\langle 1,0\rangle}_{Y}$ is a counterfactual distribution, in the sense that we don’t observe data sampled from it, and so we cannot estimate $\theta^{\langle 1,0\rangle}$ directly as a sample average. It is still possible, however, to identify $\theta^{\langle 1,0\rangle}$, as shown in the following lemma:
@@
The intuition for Lemma 2.1 is captured graphically in Figure 2. Equation (REG) gives identification by regression. We learn the dependence between $Y$ and $X$ in sample 0 through a (non-parametric) regression, and then average that regression function over the $X$ in sample 1. This is essentially a non-parametric generalization of the Oaxaca-Blinder decomposition (Oaxaca, 1973; Blinder, 1973). It also appears in the causality literature as part of the mediation formula (see Appendix B and Pearl, 2009, §4.5). Equation (REW) gives identification by re-weighting. We average $Y$ in sample 0, but we give more weight to observations whose $X$ is more likely to come from sample 1. Formally, the weights $\alpha(X)$ are Radon-Nykodim (RN) derivatives (Billingsley, 1995). When $X$ has a density or a probability mass function, these are simply the ratio of densities or probability mass functions between the two samples, respectively. The re-weighting idea has been used in the literature on covariate shift problems (Shimodaira, 2000; Bickel et al., 2009), but we are not aware of causal change attribution methods that leverage this insight.
[…]
Figure 2: Visual intuition for regression and re-weighting.
@@
Equation (DR) is doubly robust in the following sense: it still identifies the parameter of interest even if one of the regression function or the weights is misspecified. The term $\mathrm{E}_{(0)}[a(X)\{Y-g(X)\}]$ is a “debiasing” term, as in the double/debiased machine learning literature (Chernozhukov et al., 2018a), consisting of an average of the non-parametric regression error $Y-g(X)$ weighted by $a(X)$. If the regression function is correctly specified, this average will be zero regardless of the weights (by the Law of Iterated Expectations). On the other hand, if the regression function is not correctly specified but the weights are, the second term will account and correct for the misspecification of $g(X)$. We refer the reader to the proof in Appendix A for details.
@@
One of the regularity conditions (A.1) imposes that the support of $P^{(0)}_{X}$ includes the support of $P^{(1)}_{X}$. From a technical perspective, this assumption guarantees that the RN derivative $\alpha(X)$ exists. In this remark, we give a more intuitive explanation. The regression strategy (REG) requires estimating a regression of $h(Y)$ on $X$ in sample 0, and then obtaining the fitted values in sample 1. Without overlap, we would be extrapolating. In general, we want to avoid this (unless we have a credibly good parametric model for the regression). Similarly, the re-weighting strategy (REW) breaks down without overlap, because some regions of $X$ values that have positive probability in sample 1 are never observed in sample 0. ∎
@@ 2.3 Identification
Let $g_{k}(\bar{\bm{X}}_{k})$, $a_{k}(\bar{\bm{X}}_{k})$ be any functions such that $\mathrm{E}_{(c_{k+1})}[g_{k}(\bar{\bm{X}}_{k})^{2}]<\infty$, $\mathrm{E}_{(c_{k+1})}[a_{k}(\bar{\bm{X}}_{k})^{2}]<\infty$ for $k=1,\ldots,K$. Consider the following estimating equation:
$\mathrm{E}_{(c_{1})}[g_{1}(X_{1})]+\sum_{k=1}^{K}\mathrm{E}_{(c_{k+1})}[a_{k}(\bar{\bm{X}}_{k})\{g_{k+1}(\bar{\bm{X}}_{k+1})-g_{k}(\bar{\bm{X}}_{k})\}].$ (MR)
@@
$\gamma_{k}(\bar{\bm{X}}_{k}):=\mathrm{E}_{(c_{k+1})}[g_{k+1}(\bar{\bm{X}}_{k+1})\mid\bar{\bm{X}}_{k}],\quad\alpha_{k}(\bar{\bm{X}}_{k}):=\prod_{j=1}^{k}\frac{\mathrm{d}P^{(c_{j})}_{X_{j}\mid\bar{\bm{X}}_{j-1}}}{\mathrm{d}P^{(c_{k+1})}_{X_{j}\mid\bar{\bm{X}}_{j-1}}}(X_{j}\mid\bar{\bm{X}}_{j-1}).$
Under regularity conditions given in the appendix, (MR) is equal to $\theta^{\bm{c}}$ if, for every $k=1,\ldots,K$, either $g_{k}(\bar{\bm{X}}_{k})=\gamma_{k}(\bar{\bm{X}}_{k})$ or $a_{k}(\bar{\bm{X}}_{k})=\alpha_{k}(\bar{\bm{X}}_{k})$, but not necessarily both.
The intuition for this result is similar to Lemma 2.2. The parameter of interest can be estimated by regression, but now we need to use sequentially nested regressions $\gamma_{k}(\bar{\bm{X}}_{k})$ (i.e. regressions where the outcome itself is a regression function). This fixes the desired distribution for each causal mechanism. Since we are estimating $K$ regression functions, we require $K$ debiasing terms, which appear additively. Each of these debiasing terms includes a “weight” with true value $\alpha_{k}(\bar{\bm{X}}_{k})$, which is a product of the RN derivatives for the distributions that are shifted with respect to $c_{k+1}$. Including these $K$ debiasing terms makes equation (MR) multiply robust: for each causal mechanism, only the corresponding regression function or the corresponding weights need to be correctly specified, but not necessarily both. We also note that one of our conditions for multiple robustness is that $g_{k}(\bar{\bm{X}}_{k})=\gamma_{k}(\bar{\bm{X}}_{k})$ where $\gamma_{k}(\bar{\bm{X}}_{k}):=\mathrm{E}_{(c_{k+1})}[g_{k+1}(\bar{\bm{X}}_{k+1})\mid\bar{\bm{X}}_{k}]$. This is weaker than setting $\gamma_{k}(\bar{\bm{X}}_{k})=\mathrm{E}_{(c_{k+1})}[\gamma_{k+1}(\bar{\bm{X}}_{k+1})\mid\bar{\bm{X}}_{k}]$, because we do not require that $g_{k+1}$ is correctly specified, as long as $a_{k+1}(\bar{\bm{X}}_{k+1})=\alpha_{k+1}(\bar{\bm{X}}_{k+1})$.
@@
Suppose $K=2$ and the causal Markov factorization (F) is
$P_{Y\mid X_{1},X_{2}}P_{X_{2}\mid X_{1}}P_{X_{1}},$
as captured by the DAG in Figure 3. Suppose we are interested in estimating $\theta^{\bm{c}}$ for $\bm{c}=\langle 0,1,0\rangle$. The multiply-robust estimating equation in Theorem 2.4 for this case is:
$\mathrm{E}_{(0)}[g_{1}(X_{1})]+\mathrm{E}_{(1)}[a_{1}(X_{1})\{g_{2}(X_{1},X_{2})-g_{1}(X_{1})\}]+\mathrm{E}_{(0)}[a_{2}(X_{1},X_{2})\{h(Y)-g_{2}(X_{1},X_{2})\}],$
@@
Figure 3: DAG for 2.5.
@@
The general result of Theorem 2.4 implies that, when there are $K$ explanatory variables, we can identify $\theta^{\bm{c}}$ for any $\bm{c}$ with a multiply-robust estimating equation that depends on $2K$ unknown functions ($K$ regression functions and $K$ weights) that need to be estimated. For certain change vectors $\bm{c}$, however, it is possible to simplify the identifying equation to have fewer unknown functions while maintaining multiple robustness. In Appendix C, we discuss two settings where simplifications may occur. The first one is when $c_{k}=c_{k+1}$ for some $k$ (i.e., when two consecutive regressions are computed with respect to the same probability distribution). The second one is when $X_{k}$ and $X_{k+1}$ are (conditionally) independent. In particular, when all the explanatory variables are independent, for any given $\bm{c}$ we only need to estimate one regression and one RN derivative, regardless of $K$. ∎

### v4

@@ Abstract
Comparing two samples of data, we observe a change in the distribution of an outcome variable. In the presence of multiple explanatory variables, how much of the change can be explained by each possible cause? We develop a new estimation strategy that, given a causal model, combines regression and re-weighting methods to quantify the contribution of each causal mechanism. Our proposed methodology is multiply robust, meaning that it still recovers the target parameter under partial misspecification. We prove that our estimator is consistent and asymptotically normal. Moreover, it can be incorporated into existing frameworks for causal attribution, such as Shapley values, which will inherit the consistency and large-sample distribution properties. Our method demonstrates excellent performance in Monte Carlo simulations, and we show its usefulness in an empirical application. Our method is implemented as part of the Python library DoWhy (Sharma & Kiciman, 2020; Blöbaum et al., 2022).
Machine Learning, ICML
@@ 1 Introduction
Although other recent algorithms have been proposed to estimate Shapley Values, a majority of these algorithms are not multiply-robust, and generally they are based on regression only (learning the distribution of the outcome given the explanatory variables). [^1] The only exception that we are aware of is Jung et al. (2022). However, their notion of $do$-Shapley values decomposes the effect of fixing a value for the explanatory variables, rather than a distribution for each causal mechanism, as we do in this paper. They consider only discrete-valued explanatory variables, and their approach does not seem easily generalizable beyond that. As such, they will not perform well unless we have a high-quality estimator of the regression function. In high-dimensional or non-parametric settings, estimators for those objects will typically exhibit slow convergence rates, rendering traditional normal-based large sample inference (standard errors, confidence intervals or $p$-values) invalid. We provide the first formal analysis of the asymptotic properties of an estimator in the causal change attribution setting of Budhathoki et al. (2021), which allows us to perform valid inference in large samples (standard errors, confidence intervals and tests).
Causal change attribution is related to the classical problem of treatment effect estimation (Pearl, 2009; Imbens & Rubin, 2015; Peters et al., 2017) in the need to evaluate counterfactual distributions, but there are some fundamental differences. Within the treatment effects literature, the closest problem to ours is that of causal mediation. Tchetgen-Tchetgen & Shpitser (2012) gave multiple-robustness results for mediation analysis with a single mediator (explanatory variable). The case with multiple mediators and an arbitrary DAG is substantially more challenging. Although causal mediation with multiple mediators has been studied before (e.g., Daniel et al., 2015), we provide the first multiply-robust estimator in this setting. We also believe to be the first to explicitly build a connection between the causal change attribution problem and the causal mediation literature.
@@ 2.1 Setting
We want to draw a clear distinction between attribution and a related problem that has received a much larger attention in the literature, Average Treatment Effect (ATE) estimation under unconfoundedness. We depict the difference between the causal models underlying these two problems in Figure 1. In an ATE setting, $\bm{X}$ typically represents pre-treatment covariates, that affect both the outcome $Y$ and the propensity to receive a binary treatment $T$. To compute the causal effect of $T$ we need to control for those covariates, that is, compare units with similar values of $\bm{X}$. In contrast, causal change attribution acknowledges that both the distributions of $\bm{X}$ and of $Y$ given $\bm{X}$ vary depending on whether $T=0$ or $T=1$. The goal of causal change attribution is to quantify how much of the effect of $T$ on $Y$ goes through (is mediated by) each causal mechanism in this distribution. Under an unconfoundedness assumption about the assignment of $T$, the problem of causal attribution can be thought of as decomposing the ATE of $T$ on $Y$ into the Natural Direct Effect and the Natural Indirect Effect corresponding to each causal mechanism (Pearl, 2009, §4.5; Daniel et al., 2015). We describe this in more detail, including the corresponding structural equation model and potential outcomes, in Appendix B.
We note, however, that our definition of attribution does not require $T$ to be a treatment that can be administered in the “interventional” sense. As discussed, $T$ could also be an indicator for group membership (e.g., female or male), time period (e.g., before and after a certain date), or geographical location (e.g., East vs. West Coast of the US). Under which conditions can this change attribution be interpreted as causal? We will assume that the researcher knows a causal ordering of the underlying true causal graph for $(\bm{X},Y)$. This will allow us to quantify the contribution of each causal mechanism, rather than changes in the marginal distributions of explanatory variables. In practice, researchers can use a combination of causal discovery methods (e.g., the conditional independence test of Zhang et al., 2011) and domain knowledge to obtain the DAG. Finally, we also assume that there is no unobserved variable $U$ such that $T\rightarrow U$, $U\rightarrow\bm{X}$ and $U\rightarrow Y$. We discuss ways to relax this assumption in future research in Section 4.
@@ 2.2 Preliminary Example
$\theta^{\langle 1,0\rangle}=\int y\mathrm{d}P^{\langle 1,0\rangle}_{Y}(y),\quad\text{for}\quad P^{\langle 1,0\rangle}_{Y}=\int P_{Y\mid X}^{(0)}\mathrm{d}P_{X}^{(1)}.$
The fundamental challenge is that $P^{\langle 1,0\rangle}_{Y}$ is a counterfactual distribution (Pearl, 2009, §4.5), in the sense that we don’t observe data sampled from it, and so we cannot estimate $\theta^{\langle 1,0\rangle}$ directly as a sample average. It is still possible, however, to identify $\theta^{\langle 1,0\rangle}$, as shown in the following lemma:
@@
We show the intuition for Lemma 2.1 graphically in Figure 2. Equation (REG) gives identification by regression. We learn the dependence between $Y$ and $X$ in sample 0 through a (non-parametric) regression, and then average that regression function over the $X$ in sample 1. This is essentially a non-parametric generalization of the Oaxaca-Blinder decomposition (Oaxaca, 1973; Blinder, 1973). It also appears in the causality literature as part of the mediation formula (see Appendix B and Pearl, 2009, §4.5). Equation (REW) gives identification by re-weighting. We average $Y$ in sample 0, but we give more weight to observations whose $X$ is more likely to come from sample 1. Formally, the weights $\alpha(X)$ are Radon-Nykodim (RN) derivatives (Billingsley, 1995). When $X$ has a density or a probability mass function, these are simply the ratio of densities or probability mass functions between the two samples, respectively. The re-weighting idea has been used in the literature on covariate shift problems (Shimodaira, 2000; Bickel et al., 2009), but we are not aware of causal change attribution methods that leverage this insight.
Remark 2.2 (Relation to Mediation).
Causal Change Attribution is closely related to decomposing the total effect of an intervention that sets $T=1$ into the Natural Direct Effect and the Natural Indirect Effect of Pearl (2009, §4.5). We discuss this connection in detail in Appendix B. ∎
@@
Equation (DR) is doubly robust in the following sense: it still identifies the parameter of interest even if one of the regression function or the weights is misspecified. The term $\mathrm{E}_{(0)}[a(X)e(X,Y)]$ is a “debiasing” term, as in the double/debiased machine learning literature (Chernozhukov et al., 2018a), consisting of an average of the non-parametric regression error $e(Y,X)$ weighted by $a(X)$. If the regression function is correctly specified, this average will be zero regardless of the weights (by the Law of Iterated Expectations). On the other hand, if the regression function is not correctly specified but the weights are, the second term will account and correct for the misspecification of $g(X)$. We refer the reader to the proof in Appendix A for details.
@@
One of the regularity conditions (A.1) imposes that the support of $P^{(0)}_{X}$ includes the support of $P^{(1)}_{X}$. From a technical perspective, this assumption guarantees that the RN derivative $\alpha(X)$ exists. In this remark, we give a more intuitive explanation. The regression strategy (REG) requires estimating a regression of $h(Y)$ on $X$ in sample 0, and then obtaining the fitted values in sample 1. Without overlap, we would be extrapolating. In general, we want to avoid this (unless we have a credibly good parametric model for the regression). Similarly, the re-weighting strategy (REW) breaks down without overlap, because some regions of $X$ values that have positive probability in sample 1 are never observed in sample 0. For our asymptotic results in Section 2.5 we will impose a stronger form of overlap (A.3), which is analogous to the overlap assumption in ATE estimation under unconfoundedness. ∎
@@ 2.3 Identification
Let $g_{k}(\bar{\bm{X}}_{k})$, $a_{k}(\bar{\bm{X}}_{k})$ be any candidate functions such that $\mathrm{E}_{(c_{k+1})}[g_{k}(\bar{\bm{X}}_{k})^{2}]<\infty$, $\mathrm{E}_{(c_{k+1})}[a_{k}(\bar{\bm{X}}_{k})^{2}]<\infty$ for $k=1,\ldots,K$. Consider the following estimating equation:
$\mathrm{E}_{(c_{1})}[g_{1}(X_{1})]+\sum_{k=1}^{K}\mathrm{E}_{(c_{k+1})}[a_{k}(\bar{\bm{X}}_{k})e_{k}(\bar{\bm{X}}_{k+1})],$ (1)
where $e_{k}(\bar{\bm{X}}_{k+1}):=g_{k+1}(\bar{\bm{X}}_{k+1})-g_{k}(\bar{\bm{X}}_{k})$ for $k=1,\ldots,K$.
@@
$\gamma_{k}(\bar{\bm{X}}_{k})$ $:=\mathrm{E}_{(c_{k+1})}[g_{k+1}(\bar{\bm{X}}_{k+1})\mid\bar{\bm{X}}_{k}],$
$\alpha_{k}(\bar{\bm{X}}_{k})$ $:=\prod_{j=1}^{k}\frac{\mathrm{d}P^{(c_{j})}_{X_{j}\mid\bar{\bm{X}}_{j-1}}}{\mathrm{d}P^{(c_{k+1})}_{X_{j}\mid\bar{\bm{X}}_{j-1}}}(X_{j}\mid\bar{\bm{X}}_{j-1}).$
Under regularity conditions given in the appendix, (1) is equal to $\theta^{\bm{c}}$ if, for every $k=1,\ldots,K$, either $g_{k}(\bar{\bm{X}}_{k})=\gamma_{k}(\bar{\bm{X}}_{k})$ or $a_{k}(\bar{\bm{X}}_{k})=\alpha_{k}(\bar{\bm{X}}_{k})$, but not necessarily both.
The intuition for this result is similar to Lemma 2.3. The parameter of interest can be estimated by regression, but now we need to use sequentially nested regressions $\gamma_{k}(\bar{\bm{X}}_{k})$ (i.e. regressions where the outcome itself is a regression function). This fixes the desired distribution for each causal mechanism. Since we are estimating $K$ regression functions, we require $K$ debiasing terms, which appear additively. Each of these debiasing terms includes a “weight” with true value $\alpha_{k}(\bar{\bm{X}}_{k})$, which is a product of the RN derivatives for the distributions that are shifted with respect to $c_{k+1}$. Including these $K$ debiasing terms makes equation (1) multiply robust: for each causal mechanism, only the corresponding regression function or the corresponding weights need to be correctly specified, but not necessarily both. We also note that one of our conditions for multiple robustness is that $g_{k}(\bar{\bm{X}}_{k})=\gamma_{k}(\bar{\bm{X}}_{k})$ where $\gamma_{k}(\bar{\bm{X}}_{k}):=\mathrm{E}_{(c_{k+1})}[g_{k+1}(\bar{\bm{X}}_{k+1})\mid\bar{\bm{X}}_{k}]$. This is weaker than setting $\gamma_{k}(\bar{\bm{X}}_{k})=\mathrm{E}_{(c_{k+1})}[\gamma_{k+1}(\bar{\bm{X}}_{k+1})\mid\bar{\bm{X}}_{k}]$, because we do not require that $g_{k+1}$ is correctly specified, as long as $a_{k+1}(\bar{\bm{X}}_{k+1})=\alpha_{k+1}(\bar{\bm{X}}_{k+1})$.
@@
Suppose $K=2$ and the causal Markov factorization (F) is $P_{Y\mid X_{1},X_{2}}P_{X_{2}\mid X_{1}}P_{X_{1}}$, as captured by the DAG in Figure 3. Suppose we are interested in estimating $\theta^{\bm{c}}$ for $\bm{c}=\langle 0,1,0\rangle$. The multiply-robust estimating equation in Theorem 2.5 for this case is:
$\mathrm{E}_{(0)}[g_{1}(X_{1})]\\ +\mathrm{E}_{(1)}[a_{1}(X_{1})\{g_{2}(X_{1},X_{2})-g_{1}(X_{1})\}]\\ +\mathrm{E}_{(0)}[a_{2}(X_{1},X_{2})\{h(Y)-g_{2}(X_{1},X_{2})\}],$
@@
Figure 3: DAG for 2.6.
Remark 2.7 (Identification with Covariates).
As discussed in Appendix B, under the assumption that $T$ is a randomly assigned treatment, the causal change attribution problem has an interpretation in terms of mediation Pearl (2009, §4.5). Sometimes, random assignment of $T$ may be implausible, but the researcher may have access to a set of pre-treatment covariates $\bm{W}$ such that an unconfoundedness assumption holds conditional on $\bm{W}$. It is still possible to get multiple robustness in that case, with two adjustments: (i) the regression and weights need to control for $\bm{W}$, and (ii) an additional debiasing term is required, to account for the adjustment on covariates. This is explained in Appendix C. ∎
@@
The general result of Theorem 2.5 implies that, when there are $K$ explanatory variables, we can identify $\theta^{\bm{c}}$ for any $\bm{c}$ with a multiply-robust estimating equation that depends on $2K$ unknown functions ($K$ regression functions and $K$ weights) that need to be estimated. For certain change vectors $\bm{c}$, however, it is possible to simplify the identifying equation to have fewer unknown functions while maintaining multiple robustness. In Appendix D, we discuss two settings where simplifications may occur. The first one is when $c_{k}=c_{k+1}$ for some $k$ (i.e., when two consecutive regressions are computed with respect to the same probability distribution). The second one is when $X_{k}$ and $X_{k+1}$ are (conditionally) independent. In particular, when all the explanatory variables are independent, for any given $\bm{c}$ we only need to estimate one regression and one RN derivative, regardless of $K$. ∎